OpenStudy (anonymous):

Prove, in complete sentences, whether the left expression is equal to the right expression and discuss which property applies. 2(x − 3) <-----> 2x − 6

OpenStudy (anonymous):

As mentioned before by many helpful people, we take a look at the distributive property :) As can be seen, if we do multiply 2 by x and then 2 by -3, we do in fact get the equation on the right! In other words, 2(x-3) = 2*x - 2*3 = 2x - 6 Hope this helped! Have a nice day :)

OpenStudy (anonymous):

Thank you!!!

OpenStudy (anonymous):

You're welcome :) Sorry that I did not exactly guide you through. This is a question that you will need to practice more on to get the hang of it. I will try to guide instead of just showing the answer in the future, though. Have a great day! :)